OpenStudy (anonymous):

Please help. What is the greatest positive value for y in the system: x^2 + y^2 = 5 & xy = 2

ganeshie8 (ganeshie8):

u wana use calculus/graphing ?

OpenStudy (anonymous):

you can also use lagrange ,if you are familiar with it

OpenStudy (anonymous):

does it matter wich one you maximise and which one is your constraint

OpenStudy (anonymous):

we can say y=2/x put in the eq

OpenStudy (anonymous):

u get a new eq solve for it u get the result

OpenStudy (anonymous):

i think the geometric approach is quite obvious,the functions are are already given,the question wud be what is the highest value of y when a circle of radius sqrt{5} intersect with hyperbola,which is pbviously the intersection

OpenStudy (raden):

just an alternative x = 2/y (2/y)^2 + y^2 = 5 4/y^2 + y^2 = 5 4 + y^4 = 5y^2 y^4 - 5y^2 + 4 = 0 (y^2 - 1)(y^2 - 4) = 0 (y+1)(y-1)(y+2)(y-2) = 0 make each factor equals zero, then solve for y

OpenStudy (anonymous):



OpenStudy (anonymous):

So what would be the exact answer? :O

OpenStudy (anonymous):

@RadEn is finding the intersection ,so we really dont need maximising ,just point P,as it is evidently the highest value of y

ganeshie8 (ganeshie8):

yes just solving the equations wil do lol

OpenStudy (anonymous):

got it. Highest value is 2. SS is (1,-1,2,-2) Thanks!

OpenStudy (anonymous):

yes, calculus wud work if some function \(f(x,y)=x^2+y^2\) with \(xy=2\) as constraint 2 is the highest value,

OpenStudy (anonymous):

thanks!